# Transverse spin effects and light-quark dipole moments at lepton colliders

### v1

@@ Abstract
We thank C. Cocuzza, A. Prokudin, and N. Sato for sharing the dihadron fragmentation function code and helpful discussions. Xin-Kai Wen is supported in part by the National Science Foundation of China under Grants No. 11725520, No. 11675002 and No. 12235001. Bin Yan is supported in part by the National Science Foundation of China under Grant No. 12422506, the IHEP under Grant No. E25153U1 and CAS under Grant No. E429A6M1. Zhite Yu is supported in part by the U.S. Department of Energy (DOE) Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates Jefferson Lab, and C.-P. Yuan is supported by the U.S. National Science Foundation under Grant No. PHY-2310291.

### v2

@@ VII Acknowledgments
We thank C. Cocuzza, A. Prokudin, and N. Sato for sharing the dihadron fragmentation function code and helpful discussions. Xin-Kai Wen is supported in part by the National Natural Science Foundation of China under Grants No. 12235001 and No. 12342502. Bin Yan is supported in part by the National Natural Science Foundation of China under Grant No. 12422506, the IHEP under Grant No. E25153U1 and CAS under Grant No. E429A6M1. Zhite Yu is supported in part by the U.S. Department of Energy (DOE) Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates Jefferson Lab, and C.-P. Yuan is supported by the U.S. National Science Foundation under Grant No. PHY-2310291.